OpenStudy (anonymous):

I have moooorrreee!! So first off, you are skydiving. The air resistance slows u down at 7.2 m/s2 and you are 67.2 KG, but somehow, gravity is added to the problem too. The second one is the exact same, except air resistance is 3.29 m/s2

OpenStudy (anonymous):

[blank] m/s2 = meters per second squared, that is the acceleration

OpenStudy (sleepyhead314):

gravity's acceleration is 9.81 m/s^2

OpenStudy (anonymous):

(and work needs to be shown too, not just answer)

OpenStudy (sleepyhead314):

@jim_thompson5910

OpenStudy (anonymous):

YES, @jim_thompson5910

OpenStudy (sleepyhead314):

I am not sure of what you are looking for with the given information it is a known fact for physics problems that gravity is 9.81 m/s^2 what are you looking for?

OpenStudy (anonymous):

Newtons for both

OpenStudy (anonymous):

like it is weird, like they want newtons but when I calculated it, with gravity, i did -7.2 + 9.81 and multiplied that by the weight, but that was wrong. What was wrong there?

OpenStudy (sleepyhead314):

well you are going down gravity is pulling you down at [technically] -9.81 and air resistance is pushing up = +7.2 also, depending on the textbook/teacher/education system they might want you to use 9.8 or even 10 m/s^2 for gravity

OpenStudy (anonymous):

no my teacher wants 9.81, and so it would just be a negative of what I have?

OpenStudy (sleepyhead314):

hopefully ^_^"

OpenStudy (anonymous):

I have 175.39 N right now

jimthompson5910 (jim_thompson5910):

`So first off, you are skydiving. The air resistance slows u down at 7.2 m/s2 and you are 67.2 KG, but somehow, gravity is added to the problem too.` the problem seems incomplete?

OpenStudy (sleepyhead314):

have no fear, jim is here

OpenStudy (anonymous):

And it needs to be right, otherwise, i cant do a major grade project, which in turn fails me

OpenStudy (anonymous):

Oh and @jim_thompson5910 I need to find N

jimthompson5910 (jim_thompson5910):

They want the force pulling you down?

jimthompson5910 (jim_thompson5910):

If you can, post a screenshot of the full problem

OpenStudy (anonymous):

I don't, er, have the problem. Just the numbers, and the "slight" imagination my teacher has to make it more "fun" so that is all we have too.

jimthompson5910 (jim_thompson5910):

So like this problem, everything seems to be up in the air. That's not good

OpenStudy (anonymous):

Lol, yeah. And I attempted it, and it was wrong. (But idk if it was the #s, or just the sign)

jimthompson5910 (jim_thompson5910):

how do you know it's wrong? are you typing this into a computer system?

OpenStudy (anonymous):

No, during class, my teacher said it was wrong, both of them actually, and told me to do them for homework, and bring them tomorrow.

jimthompson5910 (jim_thompson5910):

I see

OpenStudy (anonymous):

It might be something as simple as the sign, or something as complex as all of the #s.

jimthompson5910 (jim_thompson5910):

well if there was no air resistance, then the acceleration of gravity g = 9.8 pulls you down to earth with a force of F = m*a F = m*g F = 67.2*9.8 F = 658.56 newtons

OpenStudy (anonymous):

But there is air resistance too, so UGHUGHUHGHGUGHUGHUGHUGHUGHUHGUHU

jimthompson5910 (jim_thompson5910):

`The air resistance slows u down at 7.2 m/s2` this implies that the air resistance acceleration is 9.8 - 7.2 = 2.6 m/s^2 so you're being pulled down with a force of f = m*a = 67.2*7.2 = 483.84 N the air is pushing up with a force of f = m*a = 67.2*2.6 = 174.72 N

jimthompson5910 (jim_thompson5910):

if you assign down to be negative, then the force pulling you down is -483.84 N the force pushing up is +174.72 N

OpenStudy (anonymous):

I made a mistake! Oh goodness. But the second one?

jimthompson5910 (jim_thompson5910):

so the net force is (force pulling down) + (force pushing up) = (-483.84 N) + (174.72 N) = -309.12 N

jimthompson5910 (jim_thompson5910):

which makes sense because ultimately you're going to be pulled down to earth (and not be pushed up)

OpenStudy (sleepyhead314):

wait I'm confused? shouldn't you be pulled down at 67.2*9.81 and pushed up at 67.2*7.2 ?

OpenStudy (anonymous):

That is true @jim_thompson5910

jimthompson5910 (jim_thompson5910):

`shouldn't you be pulled down at 67.2*9.81` only if there is no air resistance to slow you down if there was no air resistance, then you accelerate towards the earth at 9.81 m/s^2 but there is air resistance slowing you down (according to the problem, it's 7.2 m/s^2)

OpenStudy (sleepyhead314):

mmm ok then

OpenStudy (anonymous):

I'm back!

OpenStudy (anonymous):

Wait @jim_thompson5910 what about the second one? is it basically #1? but with changed #s?

jimthompson5910 (jim_thompson5910):

honestly I have no idea because these problems/instructions are too vague. But if you're looking for the net force (like I showed in #1), then yes you follow the same steps

OpenStudy (anonymous):

so the a is 6.52

OpenStudy (anonymous):

6.52 x 67.2 = 438.14

jimthompson5910 (jim_thompson5910):

that would be the force pulling you down

OpenStudy (anonymous):

I subtracted 9.81 - 3.29

jimthompson5910 (jim_thompson5910):

where are you getting 3.29 ?

jimthompson5910 (jim_thompson5910):

oh nvm

OpenStudy (anonymous):

so yeah!

OpenStudy (cheesecakekitten):

Someone help please: http://openstudy.com/study#/updates/5623f62de4b05fbf23f0a520

OpenStudy (chealyn98):

What's the problem